OpenStudy (anonymous):

find the slope intercept equation for the line tangent to the curve, y=f(x) when x=3, if f(3)=4 and f'(3)=-2

OpenStudy (amistre64):

they give you a point and a slope ....

OpenStudy (amistre64):

given a point (a,b), and a slope m; a line equation can be formed as: y = m(x-a)+b

OpenStudy (anonymous):

@amistre64 So I am confused how to figure out the slope and the m from the equation given?

OpenStudy (amistre64):

f' IS the slope at a given point, and the point is defined for you as x=3, f(3)=4

OpenStudy (amistre64):

given a point (a,b), and a slope m; a line equation can be formed as: (3,4) f' = -2 y = m(x-a)+b = -2(x-3)+4 expand to match the slope intercept format

OpenStudy (anonymous):

thats, y=-2x+6+4 y=-2x+10

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

@amistre64 thanks.